Ok. So i know the answer is going to be an imaginary number. The problem is this: 6 +or- the square root of (-6)^2-4(2)(9) all divided by 4. PLease help. I can not understand how to break down the problem into imaginary numbers.
Answer by lwsshak3(6456)

Ok. So i know the answer is going to be an imaginary number. The problem is this: 6 +or- the square root of (-6)^2-4(2)(9) all divided by 4. PLease help. I can not understand how to break down the problem into imaginary numbers.
**
{6±√[(-6)^2-4*2*9]}/4
{6±√[36-72]}/4
(6±√-36)/4
(6±6i)/4
6(1±i)/4
3(1±i)/2=3(1+i)/2 and 3(1-i)/2
..
note: √-36=√36*√-1=6i